# ReNF: Rethinking the Design Space of Neural Long-Term Time Series Forecasters

### v4

@@ E.1 More Ablations
Figure 12: Visualizations of the representations of different layers/sub-forecasters. (a). representations of using DO with the ETTh1 dataset. (b). representations of using BDO with the ETTh1 dataset. (c). representations of using DO with the ETTm1 dataset. (d). representations of using BDO with the ETTm1 dataset.
@@
EMA Smoothing. The results in Table 8 confirm that EMA smoothing yields substantial improvements in final forecast accuracy. This empirically validates our hypothesis from Section 2: EMA mitigates the detrimental effects of flawed early stopping, where volatile validation scores lead to the premature saving of suboptimal models. By providing a more reliable and stable training signal, EMA not only enhances performance but also establishes the robust foundation necessary to fairly evaluate our other contributions, such as the BDO paradigm.
Variants
ReNF ReNF
@@
Variants
ReNF ReNF
@@
Pre-Dropout. Our MLP architecture applies dropout to the input data by default as a regularization technique. While its impact can be subtle and dataset-dependent, we perform an ablation study for the sake of completeness. The results of this analysis on the ETTh1, ETTh2, and Electricity datasets are presented in Table 10.
Variants
ReNF ReNF
origin w/o pre-drop
Metric
MSE
MAE
MSE
MAE
ETTh1
96 0.350 0.383 0.352 0.386
192 0.385 0.408 0.385 0.410
336 0.405 0.425 0.406 0.429
720 0.422 0.449 0.428 0.454
Avg. 0.391 0.416 0.393 0.420
ETTh2
96 0.261 0.329 0.260 0.329
192 0.320 0.370 0.321 0.370
336 0.346 0.394 0.348 0.396
720 0.381 0.423 0.383 0.425
Avg. 0.327 0.379 0.328 0.380
Electricity
96 0.118 0.210 0.119 0.210
192 0.138 0.229 0.138 0.229
336 0.151 0.244 0.150 0.243
720 0.173 0.266 0.175 0.267
Avg. 0.145 0.237 0.146 0.237
Table 10: Effects of pre-drop on selected datasets.
Variants
ReNF ReNF ReNF
origin w/o factor_1 w/o factor_2
Metric
MSE
MAE
MSE
MAE
MSE
MAE
Electricity
96 0.118 0.210 0.119 0.211 0.123 0.215
192 0.138 0.229 0.138 0.229 0.145 0.236
336 0.151 0.244 0.153 0.245 0.156 0.251
720 0.173 0.266 0.179 0.271 0.177 0.269
Avg. 0.145 0.237 0.147 0.239 0.150 0.243
ETTh1
96 0.350 0.383 0.356 0.382 0.356 0.382
192 0.385 0.408 0.394 0.406 0.394 0.407
336 0.405 0.427 0.420 0.427 0.420 0.425
720 0.422 0.449 0.432 0.454 0.436 0.456
Avg. 0.391 0.417 0.400 0.417 0.402 0.418
ETTm1
96 0.270 0.325 0.273 0.327 0.275 0.327
192 0.310 0.352 0.311 0.355 0.312 0.353
336 0.343 0.373 0.343 0.376 0.342 0.374
720 0.401 0.406 0.398 0.409 0.413 0.410
Avg. 0.331 0.364 0.331 0.367 0.336 0.366
Solar
96 0.157 0.202 0.165 0.201 0.170 0.217
192 0.174 0.210 0.185 0.215 0.195 0.226
336 0.180 0.219 0.185 0.219 0.189 0.233
720 0.190 0.225 0.197 0.227 0.197 0.238
Avg. 0.176 0.214 0.183 0.216 0.188 0.229
Table 11: Ablations on the two factors of our BDO implementation. The factor_1 denotes the concatenation in input space, and factor_2 denotes the computation of the sub-forecasting losses.
E.2 Roubusteness
In Table 9, we present the error bar of ReNF in datasets with relatively small sizes or high instability. It shows that ReNF exhibits high robustness because of its simple structure, which is a favorable characteristic for industrial applications.
E.3 Further Explorations
Two Factors of BDO. We wish to clarify that our BDO paradigm is comprised of two distinct and essential mechanisms. The first is the addition of a linear head to each block, which generates an explicit sub-forecast from the intermediate representation. The second is the recursive concatenation of this sub-forecast with the input for the subsequent stage, which encourages the network to implicitly learn a post-combination function as motivated by MNFP. To disentangle their respective contributions, we conduct an ablation study on these two factors.
The results on the Electricity, ETTh1, and ETTm1 datasets, shown in Table 11, indicate that the two components are synergistic; removing either one leads to a notable degradation in performance. Overall, we find that the hierarchical supervision from the sub-forecasts plays a more significant role. This is expected, as it not only allows for repeated reuse of label information but also enforces a causal and homogeneous structure on the predictive representations across layers, an effect we visualize in Figure 12. However, a few anomalous results suggest that other unresolved factors, such as the quality of the ground truth labels, may also influence performance in certain cases.
Preference of Deep Representations. In Sec. 2.4, we introduce two variants of MLPs with a few differences. The primary distinction is the inclusion of skip-connections between representations in ReNF-$\beta$, which facilitates the learning of deeper, more complex non-linear dynamics. To justify this design choice, the following experiments demonstrate the performance degradation that occurs when a model’s complexity is mismatched with the dataset’s intrinsic characteristics.
Variants
ReNF ReNF
origin exchange
Metric
MSE
MAE
MSE
MAE
ETTh1
96 0.350 0.383 0.384 0.410
192 0.385 0.408 0.407 0.428
336 0.405 0.425 0.447 0.456
720 0.422 0.449 0.502 0.493
Avg. 0.391 0.416 0.435 0.447
ETTh2
96 0.261 0.329 0.271 0.332
192 0.320 0.370 0.340 0.376
336 0.346 0.394 0.381 0.410
720 0.381 0.423 0.415 0.437
Avg. 0.327 0.379 0.352 0.389
Electricity
96 0.118 0.210 0.125 0.217
192 0.138 0.229 0.144 0.235
336 0.151 0.244 0.160 0.252
720 0.173 0.266 0.196 0.283
Avg. 0.145 0.237 0.156 0.247
Traffic
96 0.335 0.226 0.359 0.241
192 0.356 0.239 0.377 0.249
336 0.366 0.246 0.389 0.259
720 0.402 0.267 0.426 0.284
Avg. 0.365 0.245 0.388 0.258
Table 12: Performances degrade drastically after using NFs with an improper degree of complexity. The Exchange column denotes that the version of ReNF ($\alpha$ and $\beta$) is alternated.
Appendix F Limitation and Further Discussion
The Multiple Neural Forecasting Proposition (MNFP) presented in this paper is foundational but preliminary. While it provides the core intuition for our work, a more rigorous and deeper exploration of its theoretical properties or deriving other related theorems could inspire new and promising research directions.
Specifically, under the paradigm of MNFP, the role of Neural Network (NN) in this area becomes transparent. First and perhaps most importantly, we should leverage NN’s capability to create a powerful post-combination function. While our BDO paradigm is effective, the current recursive strategy for combining sub-forecasts is not fully optimal. This is evident from the performance gap between our final forecast and the theoretical empirical bound as the number of stages increases. Developing more intricate methods to better leverage the full set of sub-forecasts could lead to substantial accuracy gains. Furthermore, the benefit of BDO is less pronounced on certain datasets, such as ETTm2. The underlying reasons for this variance warrant further investigation.
Second, we are consistently supposed to build more powerful Neural Forecasting Machines (NFMs) to approximate the expected distributions of future data, thereby reducing the bias $b$ in the MNFP C. Therefore, a comprehensive study is needed to verify the effects of our proposed techniques when applied to other advanced model architectures beyond MLPs.
Finally, this work develops the BDO paradigm specifically for the LTSF setting. A key open question is how this paradigm can be adapted for short-term forecasting tasks, which could ultimately lead to a more unified framework for time series forecasting.
@@ Appendix G Visualization of Prediction
Figure 13: Full visualization of forecasting results of ReNF. The figure shows multiple outputs of ReNF in different layers, along with the result of applying optimal post-combination.
@@
Figure 14: Full visualization of forecasting results of ReNF. The figure shows multiple outputs of ReNF in different layers, along with the result of applying optimal post-combination.
Appendix H declaration of the LLMs use
We utilized a Large Language Model (LLM) to assist in refining the language of this manuscript. The LLM was used specifically to check for grammatical correctness and to improve the clarity and flow of expressions, ensuring a professional academic tone. The sole purpose of using the LLM was to enhance the readability and comprehensibility of the paper based on our first draft. All other contents, including the core ideas, presentation logic, experimental design, and results, are entirely our own.

### v5

@@ Appendix E Further Ablations
Figure 12: Visualizations of the representations of different layers/sub-forecasters. (a). representations of using DO with the ETTh1 dataset. (b). representations of using BDO with the ETTh1 dataset. (c). representations of using DO with the ETTm1 dataset. (d). representations of using BDO with the ETTm1 dataset. This indicates that BDO tends to form predictive representations that are homogeneous and hierarchical.
E.1 The Two Principles of BDO
Method ReNF (Ours) w/o Concat w/o Supervision
MSE MAE MSE MAE MSE MAE
Electricity 96 0.118 0.210 0.119 0.211 0.123 0.215
192 0.138 0.229 0.138 0.229 0.145 0.236
336 0.151 0.244 0.153 0.245 0.156 0.251
720 0.173 0.266 0.179 0.271 0.177 0.269
Avg. 0.145 0.237 0.147 0.239 0.150 0.243
ETTh1 96 0.350 0.383 0.356 0.382 0.356 0.382
192 0.385 0.408 0.394 0.406 0.394 0.407
336 0.405 0.427 0.420 0.427 0.420 0.425
720 0.422 0.449 0.432 0.454 0.436 0.456
Avg. 0.391 0.416 0.400 0.417 0.402 0.418
ETTm1 96 0.270 0.325 0.273 0.327 0.275 0.327
192 0.310 0.352 0.311 0.355 0.312 0.353
336 0.343 0.373 0.343 0.376 0.342 0.374
720 0.401 0.406 0.398 0.409 0.413 0.410
Avg. 0.331 0.364 0.331 0.367 0.336 0.366
Solar 96 0.157 0.202 0.165 0.201 0.170 0.217
192 0.174 0.210 0.185 0.215 0.195 0.226
336 0.180 0.219 0.185 0.219 0.189 0.233
720 0.190 0.225 0.197 0.227 0.197 0.238
Avg. 0.176 0.214 0.183 0.216 0.188 0.229
Table 8: Ablation study on the core components of BDO. ’Concat’ denotes recursive concatenation in input space; ’Supervision’ denotes block-wise supervision. The best results are highlighted in bold.
(a) Ablation of BDO on M4 Datasets.
(b) Ablation of BDO on PEMS Datasets.
(c) Variation of the Layerwise Supervision Gain across PEMS Datasets
Figure 13: The ablation of two components of BDO. ’w/o sup’ removes the block-wise supervision, and ’w/o con’ denotes the ReNF removing the block-wise concatenation (thus losing the function of implicit forecast combination). We select TimeMixer wang2024timemixer and TimeMixer++ timemixer++ as the baseline for M4 and PEMS, respectively. (c) We record the variation of performance gain of block-wise supervision across different forecasting tasks. The advantage of block-wise supervision becomes clearly more significant in the setting of long-term forecasting.
We posit that our BDO paradigm derives its efficacy from two distinct and essential mechanisms: (1) block-wise Supervision, which generates explicit sub-forecasts from intermediate representations at each stage; and (2) Recursive Concatenation, which feeds these sub-forecasts back into the input of the subsequent stage, encouraging the network to implicitly learn a post-combination function. To disentangle their respective contributions across diverse forecasting tasks, we conduct ablation studies on these two factors.
Multivariate Long-Term Forecasting: Table 8 presents the ablation results on the Electricity, ETTh1, and ETTm1 datasets. The results indicate that the two components are synergistic; removing either leads to a notable degradation in performance. Overall, we observe that hierarchical supervision plays the dominant role in this setting. This is expected, as it not only facilitates the repeated reuse of label information but also enforces a causal and homogeneous structure on the predictive representations across layers—an effect visualized in Figure 12.
Short-Term Forecasting: We extend this analysis to short-term tasks using the multivariate PEMS datasets (PEMS03, PEMS08) and the univariate M4 dataset (Yearly, Quarterly, Monthly). For the M4 dataset, we adapt our strategy due to its specific characteristics (fixed, short horizons, and heterogeneous series). We set the length of each sub-forecast equal to the full forecasting horizon. This disables the short-to-long curriculum aspect of block-wise supervision, but keeps the function of implicit forecast combination. As shown in Figure 13, both components remain effective. However, in stark contrast to the long-term setting, the contribution of block-wise supervision is diminished. This aligns with our discussion in Section D.2: when the horizon is short, intermediate supervision steps may introduce auto-regressive noise and error that outweigh the information gain. To verify this hypothesis, Figure 13 illustrates the variation of performance gain provided by block-wise supervision as the forecasting horizon increases. We compare two settings on the PEMS datasets: (1) a short-term setting (forecasting 12, 24, 48 steps given 96 steps) and (2) a long-term setting (forecasting 96, 192, 336 steps given 512 steps). The results clearly show that as the task switches to long-term forecasting, the benefits of block-wise supervision become significant. This confirms that the full BDO paradigm, including its curriculum supervision, is most potent when the input and output contexts are sufficiently long.
E.2 Preference of Deep Representations.
In Sec. 2.3, we introduce two variants of MLPs with a few differences. The primary distinction is the inclusion of skip-connections between representations in ReNF-$\beta$, which facilitates the learning of deeper, more complex non-linear dynamics. To justify this design choice, the following experiments demonstrate the performance degradation that occurs when a model’s complexity is mismatched with the dataset’s intrinsic characteristics.
Dataset ReNF (Origin) ReNF (Exchange)
MSE MAE MSE MAE
ETTh1 96 0.350 0.383 0.384 0.410
192 0.385 0.408 0.407 0.428
336 0.405 0.425 0.447 0.456
720 0.422 0.449 0.502 0.493
Avg. 0.391 0.416 0.435 0.447
ETTh2 96 0.261 0.329 0.271 0.332
192 0.320 0.370 0.340 0.376
336 0.346 0.394 0.381 0.410
720 0.381 0.423 0.415 0.437
Avg. 0.327 0.379 0.352 0.389
Electricity 96 0.118 0.210 0.125 0.217
192 0.138 0.229 0.144 0.235
336 0.151 0.244 0.160 0.252
720 0.173 0.266 0.196 0.283
Avg. 0.145 0.237 0.156 0.247
Traffic 96 0.335 0.226 0.359 0.241
192 0.356 0.239 0.377 0.249
336 0.366 0.246 0.389 0.259
720 0.402 0.267 0.426 0.284
Avg. 0.365 0.245 0.388 0.258
Table 9: Impact of model complexity on performance. Origin: The optimal ReNF configuration found via search. Exchange: Results when the ReNF variant ($\alpha$ or $\beta$) is alternated, showing performance degradation when model complexity does not match data characteristics.
E.3 EMA Smoothing
@@
The results in Table 10 confirm that EMA smoothing yields substantial improvements in final forecast accuracy. This empirically validates our hypothesis from Section 2: EMA mitigates the detrimental effects of flawed early stopping, where volatile validation scores lead to the premature saving of suboptimal models. By providing a more reliable and stable training signal, EMA not only enhances performance but also establishes the robust foundation necessary to fairly evaluate our other contributions, such as the BDO paradigm.
Variants ReNF ReNF
@@
Variants ReNF ReNF
@@
E.4 Pre-Dropout.
Variants
ETTh1 ETTh2 Electricity
Metric 96 192 336 720 Avg. 96 192 336 720 Avg. 96 192 336 720 Avg.
ReNF MSE 0.350 0.385 0.405 0.422 0.391 0.261 0.320 0.346 0.381 0.327 0.118 0.138 0.151 0.173 0.145
MAE 0.383 0.408 0.425 0.449 0.416 0.329 0.370 0.394 0.423 0.379 0.210 0.229 0.244 0.266 0.237
ReNF w/o pre-drop MSE 0.352 0.385 0.406 0.428 0.393 0.260 0.321 0.348 0.383 0.328 0.119 0.138 0.150 0.175 0.146
MAE 0.386 0.410 0.429 0.454 0.420 0.329 0.370 0.396 0.425 0.380 0.210 0.229 0.243 0.267 0.237
Table 11: Effects of pre-drop.
Our MLP architecture applies dropout to the input data by default as a regularization technique. While its impact can be subtle and dataset-dependent, we perform an ablation study for the sake of completeness. The results of this analysis on the ETTh1, ETTh2, and Electricity datasets are presented in Table 11.
Appendix F Roubusteness
In Table 12, we present the error bar of ReNF in datasets with relatively small sizes or high instability. It shows that ReNF exhibits high robustness because of its simple structure and inherent variance reduction, which is a favorable characteristic for industrial applications.
@@ Appendix G Visualization of Prediction
In the following, we present the visualizations of multivariate long-term time series forecasting using ReNF. The predicted variable in each figure is randomly selected.
@@
Appendix H Limitation and Further Discussion
The Variance Reduction Hypothesis (VRH) presented in this paper is foundational but preliminary. While it provides the core intuition for the multiple forecasting design, a more rigorous and in-depth exploration of its theoretical properties could inspire new and promising research directions. For instance, deriving a bound beyond the independent forecasts would better connect to our proposal, identifying more intricate improvements to the current paradigm.
Specifically, under the paradigm of MNFP, the role of Neural Network (NN) in this area becomes transparent. First and perhaps most importantly, we should leverage NN’s capability to create a powerful post-combination function. While our BDO paradigm is effective, the current recursive strategy for combining sub-forecasts is not fully optimal. This is evident from the performance gap between our final forecast and the theoretical empirical bound as the number of stages increases. The core is that while BDO has created the skeleton of sequential causality at the level of network structure, methods for explicitly enforcing the specific causality in this paradigm are required. Developing more intricate methods to better leverage the full set of sub-forecasts could lead to substantial accuracy gains.
Second, we are consistently supposed to build more powerful Neural Forecasting Machines (NFMs) to approximate the expected distributions of future data, thereby reducing the bias $b$ in the MNFP B.1. Therefore, a comprehensive study is needed to verify the effects of our proposed techniques when applied to other advanced model architectures beyond MLPs.
Finally, although we evaluate the effectiveness on diverse short-term forecast benchmarks, this work develops the BDO paradigm specifically for the LTSF setting. A key open question is how this paradigm can be better adapted for diverse forecasting tasks with short length, which could ultimately lead to a more unified framework for time series forecasting.